Please help me solve this problem:
A local theater company is presenting a spring musical. Each adult ticket costs $12.00, and each child ticket costs $2.00. On opening night, the theater company earns $2,278.00 from ticket sales and sells a total of 249 tickets.
If x represents the number of adult tickets and y represents the number of child tickets, then which of the following systems of linear equations can be used to determine the number of adult and child tickets that the theater company sells on opening night?
a) x+y=2278
2x+12y=249
b) x+y=2278
12x+2y=249
c) 12x+2y=2278
x+y=249
d) 2x+12y=2278
x+y=249

Answer by ikleyn(52800)

.
Please help me solve this problem:
A local theater company is presenting a spring musical. Each adult ticket costs $12.00, and each child ticket costs $2.00. On opening night, the theater company earns $2,278.00 from ticket sales and sells a total of 249 tickets.
If x represents the number of adult tickets and y represents the number of child tickets, then which of the following systems of linear equations can be used to determine the number of adult and child tickets that the theater company sells on opening night?
a) x+y=2278
2x+12y=249

b) x+y=2278
12x+2y=249

c) 12x+2y=2278   <--- this is good.
x+y=249

d) 2x+12y=2278   
x+y=249